Given that f(x)=sqrt(x+2) is a one-to-one function, find the following:
a. f(7) b. f^-1(3)
=sgrt(x+2)
=sqrt(7+2)
=sqrt(9)
=3

*I looked up "a" in the back of the book and its right and the answer for "b" should be "5" but I don't know how to solve it.
*What I don't know is how to solve for f^-1 for any equation. Do you guys have any lesson plans for f^-1 finding the inverse.

Thank you so much:)

Answer by MathTherapy(10806)


Given that f(x)=sqrt(x+2) is a one-to-one function, find the following:
a. f(7) b. f^-1(3)
=sgrt(x+2)
=sqrt(7+2)
=sqrt(9)
=3

*I looked up "a" in the back of the book and its right and the answer for "b" should be "5" but I don't know how to solve it.
*What I don't know is how to solve for f^-1 for any equation. Do you guys have any lesson plans for f^-1 finding the inverse.

Thank you so much:)

Are you sure "b" is 5? I got 7